Suppose the perfectly competitive price is given as $46 and the total cost of the firm is given by TC=14X+2X2, find

a. The profit maximizing level of output for the firm?

b. The profit of the firm?

1
Expert's answer
2021-12-16T14:38:33-0500

a. MC=P

"MC=TC'=(14X+2X^2)'=14+4X"

14+4x=46

4x=32

x=8

b. 

"profit=MR-MC=46\times8-(14+4\times8)=368-14-32=322"